OpenStudy (anonymous):

how can i solve this fraction problem? 1/6 +4/5 showing work. thanks

OpenStudy (anonymous):

(5/30)+(24/30)=(29/30)

OpenStudy (anonymous):

u gotta show how u got that

OpenStudy (anonymous):

what do you mean? 5*6=30. you need a common denominator to add fractions. for example (1/6) has to be multiplied by (5/5) so that you get (5/30) bc (4/5)*(6/6) gets you (24/30). now you have two fractions with the same denominator and add the top accordingly